OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, – 5) with a slope of −3. y = −3x − 13 y = −3x + 11 y = −3x + 13 y = −3x + 1 Is it D?

pooja195 (pooja195):

Correct :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

want to help with another

pooja195 (pooja195):

Sure ^.^

OpenStudy (anonymous):

Given the function f(x) = 2(x + 4), find x if f(x) = 20. 6 48 14 8 is it A?

pooja195 (pooja195):

Yes that is correct

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (−2, −3) with a slope of −6. y + 3 = −6(x + 2) y + 2 = 6(x + 3) y − 3 = −6(x − 2) y − 2 = 6(x − 3) Is it C?

pooja195 (pooja195):

Nope

OpenStudy (anonymous):

dang...

pooja195 (pooja195):

Remeber that 2 negatives next to each other make a positive

pooja195 (pooja195):

Try again

OpenStudy (anonymous):

B?

pooja195 (pooja195):

Nope \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Need help? Are you stuck?

OpenStudy (anonymous):

is it D?

pooja195 (pooja195):

no y1=-3 x1=-2 \[\huge~\rm~ y--3=-6(x--2) \] \[\huge~\rm~y+3=-6(x+2) \]

OpenStudy (anonymous):

ohhh i guess that makes sense